OpenStudy (blackops2luvr):

in a flower garden, there are 6 tulips for every 7 daisies. If there are 42 tulips, how many daisies are there? 55 51 47 49

OpenStudy (blackops2luvr):

@SolomonZelman

OpenStudy (blackops2luvr):

@Nurali

OpenStudy (blackops2luvr):

@ikram002p

OpenStudy (blackops2luvr):

@hartnn

OpenStudy (anonymous):

6/7 = 42/x (6 tulips to 7 daises = 42 tulips to x daisies) cross multiply (6)(x) = (42)(7) 6x = 294 -- divide both sides by 6 x = 294/6 x = 49 there are 49 tulips